How fast will a satellite orbiting at 600 km (circular) above the surface of Earth be...

  1. How fast will a satellite orbiting at 600 km (circular) above the surface of Earth be traveling?
Answer #1

To calculate velocity of satellite,

V=√(G.M earth)/r

Where V is velocity of satellite

G= gravitational constant=6.67×10^-11 m^3kg^-1s^-2

M earth is mass of earth=5.972×10^24 kg

r is distance between satellite and earth core=6371+600=6971 km=6971000m

So V= √((6.67×10^-11)×(5.972×10^24))/6971000 =7907.13 m/s